OpenStudy (anonymous):

If f(x) = −2x + 1, find f(−6).

OpenStudy (bahrom7893):

just plug in -6 instead of x: -2(-6)+1 = 12 + 1 = 13

OpenStudy (bahrom7893):

MEDAL!

OpenStudy (saifoo.khan):

Insert the values simply.